A proton is placed in an electric field of intensity 500 N/C. What are the magnitude and direction of the acceleration of this proton due to this field? (m proton = 1.67 � 10-27 kg,e = 1.60 � 10-19 C)
A proton is placed in an electric field of intensity 500 N/C. What are the magnitude and direction of the acceleration of this proton due to this field? (m proton = 1.67 � 10-27 kg, e = 1.60 � 10-19 C)
4.79 � 109 m/s2 opposite to the electric field
47.9 � 1010 m/s2 opposite to the electric field
47.9 � 1010 m/s2 in the direction of the electric field
4.79 � 1010 m/s2 opposite to the electric field
4.79 � 1010 m/s2 in the direction of the electric field
Answer #1

1. a = Eq/m = 500 * 1.6*10^-19/ 1.67 *10^-27 = 4.79 *10^ 10 m/s^2


in the direction of electric field